1.Calculate the Molarity of a solution that contains 20.0 g NaCl in 0.30 L of solution.

2.Calculate the mass in grams of lithium nitrite of 120. mL of 0.500 M solution.

Describe how to prepare the solution.

3 Write the molecular, ionic and net ionic equations for the reactions in 26 and 27 between the following. If

there is no net ionic equation, write “no net ionic equation”:

potassium carbonate and sodium nitrate


Which of the following 1.00 m solutions has the highest boiling temperature?

a) Na2S

b) LiNO3

c) (NH4)3PO4

d) sucrose

Of the following solutions, ................ has the highest freezing temperature and

.................... has the lowest freezing temperature? Your answer should include two

letters in the correct order.

a) 5.0 m KCl

b) 3.0 m LiNO3

c) 2.0 m (NH4)3PO4

d) sucrose


30 .......................The freezing point depression constant, Kf of H2O = 1.86 oC/ m-1


The boiling point elevation constant, Kb of H2O = 0.52 oC oC/ m

Calculate the boiling temperature of a solution of 2.00 m AlCl3


1
Expert's answer
2022-03-18T17:22:03-0400

1) M = n(solute)/V(solution) = m(solute)/(Mr(solute)*V(solution)) = 20.0 g /(58.44 g/mol * 0.30 L) = 1.14 mol/L

2) M = n(solute)/V(solution) = m(solute)/(Mr(solute)*V(solution))

0.500 mol/L = m(LiNO2)/(52.95 g/mol * 0.120 L)

m(LiNO2) = 52.95 g/mol * 0.120 L * 0.500 mol/L = 3.18 g

3) K2CO3 + 2NaNO3 = Na2CO3 + 2KNO3

2K+ + CO32- + 2Na+ + 2NO3- = 2Na+ + CO32- + 2K++ 2NO3-

No net ionic equation

4) ΔTb = Kb*m*i

Kb = molal boiling point constant (0.52 °C/m for aqueous solutions)

m = 1.00 m

Na2S = 2Na+ + S2-; i = 2+1 = 3

LiNO3 = Li+ + NO3-; i = 1+1 = 2

(NH4)3PO4 = 3NH4+ + PO43-; i = 3+1 = 4

For sucrose i = 1

(NH4)3PO4 has the highest boiling temperature

5) ΔTf = Kf*m*i

Kf = molal freezing point constant (1.86 °C/m for aqueous solutions)

a) KCl = K+ + Cl-; i = 1+1 = 2

ΔTf = 1.86 °C/m * 5.0 m * 2 = 18.6°C

b) LiNO3 = Li+ + NO3-; i = 1+1 = 2

ΔTf = 1.86 °C/m * 3.0 m * 2 = 11.16°C

c) (NH4)3PO4 = 3NH4+ + PO43-; i = 3+1 = 4

ΔTf = 1.86 °C/m * 2.0 m * 4 = 14.88°C

d) For sucrose i = 1

Of the following solutions, d) sucrose has the highest freezing temperature and

a) KCl has the lowest freezing temperature?

6) ΔTb = Kb*m*i

Kb = molal boiling point constant (0.52 °C/m for aqueous solutions)

m = 2.00 m

AlCl3 = Al3+ + 3Cl-; i = 1+3 = 4

ΔTb = 0.52 °C/m * 2.00 m * 4 = 4.16°C

Tb = 100°C + 4.16°C = 104.16°C